Comparing three depreciation methods

Dexter Industries purchased packaging equipment on January 8 for $72,000. The equipment was expected to have a useful life of three years, or 18,000 operating hours, and a residual value of $4,500. The equipment was used for 7,600 hours during Year 1, 6,000 hours in Year 2, and 4,400 hours in Year 3.

Required:

1. Determine the amount of depreciation expense for the three years ending December 31, by (a) the straight-line method, (b) the units-of-activity method, and (c) the double-declining-balance method. Also determine the total depreciation expense for the three years by each method. Round the final answers for each year to the nearest whole dollar.

Depreciation Expense
Year Straight-Line Method Units-of-Activity Method Double-Declining-Balance Method
Year 1 $ $ $
Year 2 $ $ $
Year 3 $ $ $
Total $ $ $
Answer #1

Calculate depreciation expense

Year Straight line method Units of activity method Double decline balance method
Year 1 22500 7600*3.75 = 28500 72000*2/3 = 48000
Year 2 22500 6000*3.75 = 22500 72000*1/3*2/3 = 16000
Year 3 22500 4400*3.75 = 16500 3500
Total 67500 67500 67500

Straight line Depreciation = (72000-4500/3) = 22500

Depreciation expense per hour = (72000-4500)/18000 = 3.75

Double decline rate = 100/3*2 = 66.67%